1. (a) Show that ez 3 2 = + if and only if 2 = ( +20k) : + 2nk i for k e Z. 2
Answer #1

1) @ det e = 12 + 1 Z y = cos (1%+2kn) + i sin( 17+ 2x1) ЭСР {(4% +260) [1:2%= 630+isino) -- = (1% + 2km) =e celin KEZ conver